A manufacturer of GPS systems has an annual cost of goods sold of $7 million a year and an average inventory of $4.9 million. The GPS navigation systems manufacturer has hired a consultant to help them increase their inventory turns. They claim that by implementing a new IT system for managing their inventory, the GPS systems company will be able to increase its inventory turns to 4 times per year. What would be the average amount of inventory (in million $) with the new IT system implemented? Choose the answer closest to the correct answer. Select one: a. 0.25 b. 0.57 c. 1.75 d. 2.33 e. 19.6
Answer #1

Inventory Turnover Ratio = Cost of Goods Sold/Average Inventory

=> Inventory Turnover = 7/4.9 = 1.429

New Inventory Turnover = 4

=> COGS/Average Inventory = 4

=> Average Inventory = 7/4 = 1.75

Hence, (c) is the correct option